EDITORIAL

Terry Goodman

Welcome to a new academic year and the last issue of MJMS for 1989. We here at MJMS hope that your year has started well and wish you all the best.

In this issue you will find four interesting articles covering a wide range of topics. Any comments, additions, etc. that you might have about an article will be welcome. We would like to share these with our readers in a forthcoming issue. MJMS continues to solicit manuscripts and we would particularly encourage manuscripts from computer science as well as those focusing on various issues (i.e. – assessment, role of calculus in the curriculum, etc.)

You will find that the problems section continues to grow. We continue to encourage you to send us your problems and solutions to problems found in this section.

On page 49 you will find a Referee’s Application Form. If you would like to serve as a referee please fill out this form and return it to us. The referees for MJMS are extremely important and help to insure the quality of articles published in the journal.

Finally, let me encourage you to share MJMS with colleagues. As our number of subscribers grows, we will be able to include more articles and other features in MJMS. Wider participation in MJMS will only help us to share more ideas and hear more points of view. Thank you for your support of MJMS.